Hoodmemes:

Find for which value of x is the following equation true? 44x-7 = 4x - 1 . (A) 2 (B) 3 (C) 4

JammarWalker:

Have any clue?

Hoodmemes:

I did a little math \[4x - x - 7 = -1\]

JammarWalker:

Do you have a calculator?

Hoodmemes:

yes

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker Do you have a calculator? \(\color{#0cbb34}{\text{End of Quote}}\) If so just plug one number in at a time and see which one comes out even

boredfr:

44x−7=4x−1 Step 1: Subtract 4x from both sides. 44x−7−4x=4x−1−4x 40x−7=−1 Step 2: Add 7 to both sides. 40x−7+7=−1+7 40x=6 Step 3: Divide both sides by 40.

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @boredfr 44x−7=4x−1 Step 1: Subtract 4x from both sides. 44x−7−4x=4x−1−4x 40x−7=−1 Step 2: Add 7 to both sides. 40x−7+7=−1+7 40x=6 Step 3: Divide both sides by 40. \(\color{#0cbb34}{\text{End of Quote}}\) direct answers are not allowed

Hoodmemes:

Ok well I got as far as to this writing \[3x - 7 = -1\] \[3x = -1 + 7\] Then I did \[3x = 6\] I been tryan message my teacher for the last part but she won't respond.

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker \(\color{#0cbb34}{\text{Originally Posted by}}\) @boredfr 44x−7=4x−1 Step 1: Subtract 4x from both sides. 44x−7−4x=4x−1−4x 40x−7=−1 Step 2: Add 7 to both sides. 40x−7+7=−1+7 40x=6 Step 3: Divide both sides by 40. \(\color{#0cbb34}{\text{End of Quote}}\) direct answers are not allowed \(\color{#0cbb34}{\text{End of Quote}}\) this wasnt a direct answer bc I didnt complete the last step

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @JammarWalker \(\color{#0cbb34}{\text{Originally Posted by}}\) @boredfr 44x−7=4x−1 Step 1: Subtract 4x from both sides. 44x−7−4x=4x−1−4x 40x−7=−1 Step 2: Add 7 to both sides. 40x−7+7=−1+7 40x=6 Step 3: Divide both sides by 40. \(\color{#0cbb34}{\text{End of Quote}}\) direct answers are not allowed \(\color{#0cbb34}{\text{End of Quote}}\) thanks

darkknight:

y hood? what is 6/3

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight y hood? what is 6/3 \(\color{#0cbb34}{\text{End of Quote}}\) 2

boredfr:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight y hood? what is 6/3 \(\color{#0cbb34}{\text{End of Quote}}\) lmao

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Hoodmemes Ok well I got as far as to this writing \[3x - 7 = -1\] \[3x = -1 + 7\] Then I did \[3x = 6\] I been tryan message my teacher for the last part but she won't respond. \(\color{#0cbb34}{\text{End of Quote}}\) Good, you're on the right track.

JammarWalker:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight y hood? what is 6/3 \(\color{#0cbb34}{\text{End of Quote}}\) lol